OpenStudy (anonymous):

Solve each equation by factoring. Check your solutions. 4y^2=81 I have got a general idea how to do, but I get to (2y{-/+}9) (2y{-/+}9)

OpenStudy (photon336):

\[4y^{2} = 81 \]

OpenStudy (anonymous):

yes

OpenStudy (photon336):

step one we need to get by itself how do you think we can do this?

OpenStudy (anonymous):

Subtract the 81 and replace with a 0

OpenStudy (photon336):

watch so this actually means 4*y^{2} = 81 or 4 times y^{2} we need to do the opposite of multiplication which is division. so we need to first divide both sides by 4 \[\frac{ 4 }{ 4 }y^{2} = \frac{ 81 }{ 4 }\] \[y^{2} = \frac{ 81 }{ 4 }\]

OpenStudy (photon336):

you could do that also and I'll explain that too.

OpenStudy (anonymous):

ok ya i see what I was missing. So after that its y^2= 20.25

OpenStudy (anonymous):

Then I get the square root of that and its 4.5

OpenStudy (anonymous):

y=4.5

OpenStudy (photon336):

\[\sqrt{y^{2}} = \sqrt{\frac{ 81 }{ 4 }} = \pm~\frac{ 9 }{ 2 }\]

OpenStudy (anonymous):

well when you divide by the 4 why not just go ahead and get a decimal? else it could also be 20 1/4

OpenStudy (anonymous):

\[20 \frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

there we go. I coulnt figure out how you were doing it

OpenStudy (photon336):

I'm not sure how you got that. here is the way that I did this problem, what I was saying above.